Automobiles and trucks pollute the air with NO. At 2000.0°C, Kc for the reaction N2(g)+O2(g)⇌2NO(g) At 2000.0°C Kc = 4.100 × 10–4 and ΔH° = 180.6 kJ. What is the value of Kc at 31.00°C?


1
Expert's answer
2021-02-11T04:49:21-0500

Use Van't Hoff equation

"ln\frac{K_2}{K_1}=\frac{-\Delta H}{R}(\frac{1}{T_2}-\frac{1}{T_1})"


"T_1(K)= 2000.0 + 273= 2273 K"

"T_2(K) = 31.00+273=304.K"

"\Delta H = 180.6 kJ = 180600J"

"K_1 = 4.100\times10^{-4}"

"K_2 - ?"


"ln\frac{K_2}{4.100\times 10^{-4}}= \frac {-180600}{8.314}(\frac{1}{304}-\frac{1}{2273})"


"K_2 = 1.886\times10^{-26}"